Is –4 a solution to the equation 7x – 5 + 3x = 6 + x – 10
Answer by chitra(359)

The given expression is:

7x – 5 + 3x = 6 + x – 10

This can be written as: (Adding like terms)

10x - 5 = x - 4
==> plugging in x = -4 we get

10(-4) - 5 = -4 - 4
-40 - 5 = - 8

-45 is not equal -8
hence, -4 is not a solution